OpenStudy (anonymous):

Which numbers are equivalent to 0.32? Choose all answers that are correct. A. 3.2% B. 32% C. 8/25 D. 32/10

OpenStudy (anonymous):

stop tagging me in stuff u spammer idk u

OpenStudy (anonymous):

........I'm not......tagging you.....at all.... o-o

OpenStudy (anonymous):

aaaand I never have o.O

OpenStudy (triciaal):

0.32 is the same as 32/100 percent means per 100 so 32/100 = 32% 32/100 can be simplified 32= 4 *8 100 = 4* 25 so 32/100 = 4*8/4*25 = 8/25

OpenStudy (anonymous):

Thank you!

OpenStudy (triciaal):

you are welcome